OpenStudy (anonymous):

Arctan(-radical3) answer in radians and decimal approximation

OpenStudy (anonymous):

Try computing it this way. Let \(x=\arctan(-\sqrt3)\), so then \(\tan x=-\sqrt3\). Keep in mind the domain of one period of the tangent function, \(-\dfrac{\pi}{2}\) to \(\dfrac{\pi}{2}\). What angle \(x\) has a tangent of \(-\sqrt3\)?

OpenStudy (anonymous):

I know that there is difference between arctan and Arctan as well

OpenStudy (anonymous):

I wasn't aware of that. Care to elaborate?

OpenStudy (anonymous):

Well I'm not too sure but in my book it says " that the special use of the capital letter distinguishes inverse functions from corresponding inverse relations"

OpenStudy (anonymous):

\[\text{ArcTan}\left[-\sqrt{3}\right]=-\frac{\pi }{3}=-60{}^{\circ}=120{}^{\circ} \]

OpenStudy (anonymous):

ohh so its the same as arctan? ok :)

OpenStudy (anonymous):

because its funny that I would be asked to solve for arctan(-radical 3) and Arctan(-radical3) and the same for other numbers with sin and cos inverses just to get all the same answer twice

OpenStudy (anonymous):

can anyone confirm that they are both the same answers? because it is kinda odd

OpenStudy (anonymous):

arctan has to be typed "ArcTan" in Mathematica, a computer program, or the arctan function will not be invoked.

OpenStudy (anonymous):

ok